RC Meeting and Run Plan, Thu, January 23rd 2020

From HallCWiki
Jump to: navigation, search

RC Meeting Information

RC Daily Meetings are at 4:15 PM, 2nd floor Counting House Meeting Room

  • Bluejeans remote connection information:
 To join the Meeting:
 To join via Room System:
 Video Conferencing System: bjn.vc -or-
 Meeting ID : 861439920
 To join via phone :
 1)  Dial:
   +1.408.740.7256 (US (San Jose))
   +1.888.240.2560 (US Toll Free))
   +1.408.317.9253 (US (Primary, San Jose))
   (see all numbers - http://bluejeans.com/numbers)
 2)  Enter Conference ID : 861439920

Useful Links

General updates

Wednesday swing – Thursday owl

  • 7PM Beam back
  • 8PM – 11PM beam tuning; hoping for <300 um
    YAG [1] [2]
    IHA3H07B [3]
    IHA3H07A [4]
  • Decided to use a smaller raster (4.5 mm instead of 5 mm [5]) with this spot size (447 um x 351 um [6]) for now
  • 11:30PM – 7:30AM Production data; pol 3He, transverse polarization
    • Problem with beam line dipole power supply [7] [8]

Thursday day

  • 8AM – 11:30AM DOE tour
  • 10:30AM SHMS dipole PSU tripped [9].
    • Tried resetting and turning on, kept tripping.
    • Called Steve Lassiter and he went down to the hall around 11:30AM
    • SHMS dipole back up and stable at 1:15PM [10]
  • 12PM AFP loss [11]
  • 2PM Hall sweep finished; back to beam permit
  • 2PM Target polarization cross calibration [12]
  • 3PM Another SHMS dipole trip, shortly followed by HMS dipole, Q2, and Q3
    • Called Steve, who said to call Mike Fowler, who said they'd look into it.
    • Cryo problem? Hall A had similar problems [13]

Recent polarization measurements

 Long   34%  6PM Wed [14]
 Long   36%  9PM Wed [15]
 Trans  38% 10PM Wed [16]
 Trans  60%  4AM Thu [17]
 Long   58%  4AM Thu [18]
 Long?  58%  8AM Thu [19]
 Trans  54%  2PM Thu [20]
  • Since transverse polarization is high, seems cell has not degraded and don't need to rush to change it.
  • After measurements, if longitudinal is indeed lower than transverse, may need to open the enclosure.
  • Take data over weekend and open enclosure early next week
  • Take advantage of high transverse polarization, rotate to longitudinal polarization, then take data.

Transverse AFP loss test, Thursday 1/23

  • Loss per sweep is 1.4% [21]
  • Note that "per sweep" means "per single up or down sweep," not per combined up&down sweep.

Beam width

  • Tuning width below 300 um will require interruption to Hall B, so this needs to be scheduled in advance with their agreement. [22]
  • Clarify procedure described online for beam checkout?

Current Run Plan

  • Always take 120Hz of EDTM

Thursday swing – Tuesday day

!!                                                     !!
!!  All NMR measurement HCLOG entries should include:  !!
!!     1) A title that clearly lists polarization      !!
!!     2) Screenshots of both NMR_FieldSweep and       !!
!!        the polarization fit.                        !!
!!                                                     !!
!!  For an example see the following entry:            !!
!!     https://logbooks.jlab.org/entry/3768696       !!
!!                                                     !!
  • Polarized 3He target
  • The overall plan is to take 10 (i.e. 5+5) hours of data with transverse spin, then 5 hours with longitudinal, then 10 transverse, 5 longitudinal, 10 transverse, etc.
  • NMR measurements should be taken every 5 hours, including twice when changing polarization direction (1 before and 1 after). We want to have know polarization before and after each run.
  • To clarify:
    • Take NMR during the halfway point of transverse data acquisition
    • Before changing polarization directions, take an NMR measurement. Then change direction and wait 10 minutes to let things equilibrate. Then take another NMR measurement. Then begin taking data.

So, as a timeline:

Take 5 hours transverse
Take 5 hours transverse
Change direction, wait 10 minutes
Take 5 hours longitudinal
Change direction, wait 10 minutes
Take 5 hours transverse
Take 5 hours transverse
Change direction, wait 10 minutes
Take 5 hours longitudinal
Change direction, wait 10 minutes
etc ...

Moller & spin dance

  • Tentatively Monday morning, pending Hall A
  • Hall B strongly encourages us to push it back to the 30th (because it won't matter to them at that point and they currently have priority)


  • Open enclosure early next week, maybe Tuesday 1/28
    • Check longitudinal laser setup
    • Install the rotation stages?
    • Install EPR fiber bundle?
    • Check to see if there's any scheduled downtime that we can use for this
  • Arun tentatively getting replacement cell prepared early next week; install it in two(?) weeks?

Special Requests

  • Keep target expert schedule filled out on whiteboard

Notices in Effect for Shift Crew

  • Policy for CANS-related hall status change: Make a logbook entry when the hall goes in, or comes out of laser-controlled access.
  • Magnet policies
    • HMS Q1 must be ramped down in 50A steps in all cases
    • Do not ramp HMS Q2 to zero, only go to 10A and then switch off power.
  • Inform the RC of any access that the Target Team requires. The RC will coordinate with RadCon to make everything run smoothly.
  • The operators have been somewhat careful with our beam orbit locks. Please make sure that the orbit locks are enabled and the beam does not drift from its nominal position.
  • Ramp rate lock instructions are here and added to checklist[23]
  • Temporary Laser instructions can be found here [24]

Planned Accelerator Configuration Changes